In a crystalline solid, anions B are arranged in a cubic close packing. Cations A are equally distributed between octahedral and tetrahedral voids. If all the octahedral voids are occupied, what is the formula of the solid

Grade:

4 Answers

Sunil Kumar FP
10 years ago
no of atoms ina ccp unit cell-simila to fcc=4
therefore no of B atom present=4
no of A atom=tertrahedral void and octahedral void(8+4)=12
formula of the compound=A12B4
thanks and regards
sunil kr
askIITian faculty
Anjali
7 years ago
For A, no. of ov= 4, so no. of TV will also be 4 ( as equally distributed). So, A=8; B=4 (no. of atoms in fcc).... formula is A2B
SHIVANSH SUMAN
6 years ago
NO. OF ATOMS IN CCP =4
THEREFORE,NO. OF ATOMS IN CLOSE PACKING =B=4
WE KNOW OCTAHEDRAL VOIDS ARE EQUAL TO THE NO. OF ATOMS IN CLOSE PACKING
THEREFORE,OCTAHEDRAL VOIDS =4
AND TETRAHEDRAL VOIDS ARE 2(NO OF ATOMS IN CLOSE PACKIN),WHICH IS EQUAL TO 8
GIVEN, A ATOMS ARE EQUALLY DISTRIBUTED BETWEEN BOTH VOIDS AND A ATOMS COVER ALL OCTAHEDRAL VOIDS.THEREFORE TOTAL A ATOMS =4(OCTAHEDRAL ATOMS)+4(TETRAHEDRAL ATOMS)=8 ATOMS
THEREFORE FORMULA OF AB=A8B4 OR A2B.
NOTE THAT 4 TETRAHEDRAL VOIDS ARE NOT OCCUPIED
 
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to the problem
 
B = 4 (fcc lattice)
Since there are 8 tetrahedral and 4 octahedral voids in fcc out of which all 4 octahedral voids are filled and an equal number of tetrahedral voids too.
Hence, A = 4 + 4 = 8
Hence, Formula is A8B4
or, The empirical formula is A2B
 
Hope it helps.
Thanks and regards,
Kushagra